OpenStudy (anonymous):

NEED HELP! find g[f(3)]

OpenStudy (anonymous):

For all 5 problems in this assignment, use the following functions: f(x) = x + 5 g(x) = x2 -4x + 5 h(x) = 3x2 - 2x + 5

OpenStudy (anonymous):

Plug 3 in as this x g(x) = x2 -4x + 5

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

Where ever f(x) appears, you put in 3

OpenStudy (anonymous):

i got 17

OpenStudy (anonymous):

f[8]=27

OpenStudy (anonymous):

No...umm..do what moha does.

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

@moha_10 Can you please explain to Janessia

OpenStudy (anonymous):

okay i got over here 8*8-37 =??

OpenStudy (anonymous):

27

OpenStudy (anonymous):

okay we first sub on f and then on g

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

can u help me with this one @moha_10 its more difficult h[f(5)] - g[h(1)]

OpenStudy (anonymous):

okay [h(10)-g(6)] [15-17]=-2

OpenStudy (anonymous):

thats the whole answer all of tht?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

even tht [h(10)-g(6)] part?

OpenStudy (anonymous):

@moha_10

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ohk thnks so much!

OpenStudy (anonymous):

we sub and we got like above

OpenStudy (anonymous):

ur welcome